# A novel characterization of structures in smooth regression curves: from a viewpoint of persistent homology

### v3

@@ 5 Simulation studies
This section presents some applications of Theorem 3.1 to infer the monotonicity, convexity, and modality of smooth regression curves underlying the data. We simulate noisy observations from smooth regression curves and characterize the aforementioned local structures using the methodology proposed in Section 4. We infer local structures of the underlying regression curves over a grid of points in the region [-1, 1].
In addition, we compare the results with a graphical device called SiZer map, an existing procedure for exploring structures such as peaks and valleys in smooth regression curves proposed by [6]. The SiZer map visualizes significant features such as bumps simultaneously over both location and scale (bandwidth) by using a color map. The idea is to highlight the regions in the location-scale space where the curve significantly increases and decreases. The idea is motivated by the fact that the significant bumps will be at zero crossings of the derivative between regions of significant increase and decrease. Hence, the name “SiZer” is given for “SIgnificant ZERo crossings of derivatives”. The color scheme on the SiZer map is as follows: Blue is used to indicate regions where the curve is significantly increasing; Red is used to indicate regions where the curve is significantly decreasing; Purple is used to indicate regions where the curve cannot be concluded to be either significantly increasing or decreasing; Gray is used to indicate regions where there are not enough data points to make statements about the significance of the features. The dotted curves in SiZer maps show effective window widths for each bandwidth as intervals representing ±$\pm 2h$.
We provide a quantitative evaluation of local structures using the methodology proposed in Section 4, and compare the results with the respective SiZer maps. The statistical significance of the features in the estimated barcode is tabulated in the following manner. Let =$\mathcal{B}=\{[a_{i},b_{i}]:a_{i}<b_{i},i=1,\ldots,p\}$ be an observed barcode with p 0-dimensional features. Note that features in $\mathcal{B}$ are ordered in increasing order of persistence (length of the intervals). We tabulate the estimated slope of the regression line defined in Equation (4.2), and its p-value, for each $a_{i},b_{i}\in\mathcal{B}$. The estimated slope is denoted by $\boldsymbol{\hat{\beta}_{1}}$. In addition, we tabulate the critical points, denoted by $\boldsymbol{\hat{c}_{p}}$, corresponding to each $a_{i},b_{i}\in\mathcal{B}$ to compare the results with the SiZer maps. Recall that each $a_{i},b_{i}\in\mathcal{B}$ constitutes the set of critical values of the estimated regression function. This allows one to compare the observed structure inferred from the barcode with the respective SiZer maps. In particular, let $[a,b],a<b$ be an observed 0-dimensional feature in the barcode of the estimated derivative, where $a$ and $b$ denote the death and birth times, respectively. Further, let $a>0$, that is the observed death time is positive implying that $\hat{m}$ is an increasing function. Now, visualize the critical point corresponding to the death time $a$ on the SiZer map of $\hat{m}$. If the critical point lie on, say, blue region, indicating the $\hat{m}$ is significantly increasing, then both the conclusions coincide. Otherwise, the inferences drawn from the SiZer maps and the proposed procedure based on the barcode of the estimated derivative are contrary.
In what follows, we provide simulation results for the normal and Cauchy kernel supported on [-1, 1]. However, we compare the results with the SiZer maps only for the Gaussian kernel since the conventional SiZer is applicable only for the Gaussian kernel.
@@ 5.1 Monotonicity
We characterize the monotonicity of a smooth and bounded regression function using the 0-dimensional features of the barcode of the first derivative. The main idea is that the birth and death time corresponding to the 0-dimensional feature with the largest persistence in a barcode provide the maximum and minimum value of the function, respectively (see, Figure 1). Therefore, we can extract the 0-dimensional feature with the largest persistence from the barcode of the first derivative to characterize the monotonicity of the function. In particular, if the death time of the feature is non-negative, that is the minimum value of the first derivative is non-negative, then the function is non-decreasing. Similarly, if the birth time of the feature is non-positive, that is the maximum value of the first derivative is non-positive, then the function is non-increasing.
Consider the regression function ${m(x)=e^{x},x\in}$ [-1, 1] to investigate its monotonicity from kernel estimates of its derivative $m_{1}$. The kernel estimates of $m$ and $m_{1}$ are shown in Figure 2 where a Gaussian kernel truncated on [-1, 1] is used for the sample size $n=50$, and for three different bandwidths $h\in\{n^{-1/3},n^{-1/5},n^{-1/7}\}$. The error distribution is the normal distribution truncated on [-1, 1] with mean 0 and variance 0.1. It is evident from Figure 2 that $m$ is estimated well while its derivative $m_{1}$ is not for each of the bandwidths. The monotonicity (non-decreasingness) of $m$ is evident from the visual inspection of the estimates of $m$. However, a quantitative assessment, such as statistical hypothesis testing, is needed to ensure that the observed monotone structure is significant, and not an artifact of sampling noise.
Figure 2: Estimates of $m$ and $m_{1}$ using a Gaussian kernel truncated on [-1, 1].
We obtain barcodes for the estimated derivative, $\hat{m}_{1}$, using the procedure mentioned in Section 3, and use the proposed procedure articulated in Section 4 to infer the monotonicity of $m$ from the observed data. The observed barcodes for different bandwidths are $\boldsymbol{\{[0.1,1.4]\},\{[0.1,1]\},\{[0.2,0.8]\}}$ for the bandwidths $h=n^{-1/3},h=n^{-1/5},h=n^{-1/7}$, respectively. These barcodes are obtained with the filtration parameters $\epsilon=0.1$ and the discrete grid of integers $G=\{-g,-g+1,\ldots,g-1,g\}$, here $g=20$. It is evident from the death times of the observed features in the barcode that the function underlying the observed data is strictly increasing for each of the three bandwidths. The true and estimated barcodes are depicted in Figure 3 where the gray dotted lines intersect the time axis at the death times of the features. Note that the time axis represents the values $t_{i},i\in G$, where $t_{i}$ is defined in Equation (3.6).
Figure 3: Estimated barcodes for $n=50$ for a Gaussian kernel truncated on [-1, 1].
Next, we increase the sample size from $n=50$ to $n=100$ to further examine this observed monotone structure. The estimated barcodes for $n=100$ are $\boldsymbol{\{[0.1,1.8]\},\{[0.2,1]\},\{[0.2,0.8]\}}$ for each of the three bandwidths, respectively. These barcodes are shown in Figure 4, suggesting that the estimated function is strictly increasing.
Figure 4: Estimated barcodes for $n=100$ for Gaussian kernel truncated on [-1, 1].
We provide a quantitative assessment of this observed monotone structure using the procedure proposed in Section 4, and compare the results with the respective SiZer maps. It is evident from Table 1 that the observed death times ($a_{1}$) for each of the bandwidths and sample sizes are significant. Since the corresponding $\boldsymbol{\hat{\beta}_{1}}$ (see, Equation (4.2)) is negative and significant. Moreover, the critical points ($\boldsymbol{\hat{c}_{p}}$) corresponding to the death times of the observed features lie in the blue region on the SiZer map (see, Figure 5) implying that the observed structure is non-decreasing and significant at these points for each of the three bandwidths. This implies that the death times of the observed features are non-negative and significant, consistent with the results in Table 1. Note that green dots highlight the critical points, and the bandwidths $h=n^{-1/3},n^{-1/5},n^{-1/7}$ are depicted through black, gray, and red dotted lines, respectively, on the SiZer map (see, Figure 5).
Table 1: Significance of Monotonicity for Gaussian Kernel.
For $\boldsymbol{n=50}$ For $\boldsymbol{n=100}$
$\boldsymbol{h=n^{-1/3}}$ $\boldsymbol{h=n^{-1/5}}$ $\boldsymbol{h=n^{-1/7}}$ $\boldsymbol{h=n^{-1/3}}$ $\boldsymbol{h=n^{-1/5}}$ $\boldsymbol{h=n^{-1/7}}$
$\mathbf{a_{1}}$ $\mathbf{a_{1}}$ $\mathbf{a_{1}}$ $\mathbf{a_{1}}$ $\mathbf{a_{1}}$ $\mathbf{a_{1}}$
$\mathbf{\hat{c}_{p}}$1 -1 -1 -1 -1 -1
$\boldsymbol{\hat{\beta}_{1}}$7.5 -18.2 -26.2 -7.2 -23.5 -40.1
$\mathbf{p}$-value $\boldsymbol{<10^{-8}}$ $\boldsymbol{<10^{-8}}$ $\boldsymbol{<10^{-7}}$ $\boldsymbol{<10^{-11}}$ $\boldsymbol{<10^{-16}}$ $\boldsymbol{<10^{-16}}$
@@
Next, we demonstrate kernel estimates of $m$ and $m_{1}$ for the Cauchy kernel truncated at [-1, 1] in Figure 6. It can be seen in Figure 6 that the estimates of $m_{1}$ are too noisy for each of the bandwidths, while $m$ is quite well estimated. To investigate the monotonicity of $m$, the estimated barcodes of the derivative are shown in Figure 7. The estimated barcodes for the bandwidths $h\in\{n^{-1/3},n^{-1/5},n^{-1/7}\}$ are $\boldsymbol{\{[-0.03,1.24]\},\{[0.02,0.84]\},\{[0.04,0.73]\}}$, respectively. The monotonicity of the estimated $m$ is not evident for $h=n^{-1/3}$ as the death time in the barcode of the derivative is negative. Furthermore, the observed non-monotone structure is significant for $h=n^{-1/3}$ as it can be seen in Table 2. On the other hand, for the bandwidths $h=n^{-1/5}$ and $h=n^{-1/7}$, the observed structure of the estimated $m$ is significantly non-decreasing. Since the death times in the observed barcodes are non-negative and significant from Table 2.
Now, we increase the sample size from $n=50$ to $n=100$ to further examine the monotonicity of $m$. The observed barcodes of the derivative for the three bandwidths are $\boldsymbol{\{[-0.08,1.50],[0,0.94],[0.04,0.75]\}}$. Again, it is evident that the observed structure is significantly non-monotone for $h=n^{-1/3}$, and significantly non-decreasing for $h=n^{-1/5}$ and $h=n^{-1/7}$. The significance of the death times ($a_{1}$) in the observed barcode of the derivative is articulated in Table 2 for different bandwidths and sample sizes.
Figure 6: Kernel estimates using the Cauchy kernel truncated on [-1, 1].
Figure 7: Estimated barcodes for $n=50$ the Cauchy kernel truncated on [-1, 1].
Figure 8: Estimated barcodes for $n=100$ Cauchy kernel truncated on [-1, 1].
Table 2: Significance of Monotonicity for Cauchy Kernel.
For $\boldsymbol{n=50}$ For $\boldsymbol{n=100}$
$\boldsymbol{h=n^{-1/3}}$ $\boldsymbol{h=n^{-1/5}}$ $\boldsymbol{h=n^{-1/7}}$ $\boldsymbol{h=n^{-1/3}}$ $\boldsymbol{h=n^{-1/5}}$ $\boldsymbol{h=n^{-1/7}}$
$\mathbf{a_{1}}$ $\mathbf{a_{1}}$ $\mathbf{a_{1}}$ $\mathbf{a_{1}}$ $\mathbf{a_{1}}$ $\mathbf{a_{1}}$
$\mathbf{\hat{c}_{p}}$1 -1 -1 -1 -1 -1
$\boldsymbol{\hat{\beta}_{1}}$8.8 -21.6 -30.2 -8.8 -28.4 -45.7
$\mathbf{p}$-value $\boldsymbol{<10^{-8}}$ $\boldsymbol{<10^{-10}}$ $\boldsymbol{<10^{-9}}$ $\boldsymbol{<10^{-11}}$ $\boldsymbol{<10^{-16}}$ $\boldsymbol{<10^{-16}}$
@@ 5.2 Convexity
Note that to characterize the monotone structure of $m_{1}$ quantitatively, we need to consider the barcodes of the second derivative of $m$. However, qualitatively, the monotonicity of $m_{1}$ can be captured by comparing the barcodes of $m_{1}$ and -$m_{1}$, as also suggested by one of the reviewers. The rationale behind the argument is that for a monotone function $f$, the super-level set $D_{t}=\{x:f(x)\geq t\}$ of $f$ has only a single connected component, for every $t\in\mathbb{R}$. However, this is not a necessary condition since if $f$ is a concave shape function, then also $D_{t}$ will have only a single connected component for every $t\in\mathbb{R}$ (see, Figure 19). Therefore, to rule out such situations we have to make use the property of monotone functions that if $f$ is monotone increasing (decreasing) on an interval $[a,b]$ (here we take a = -1, b =1), then -$f$ will be monotone decreasing (increasing) on $[a,b]$. This property allows us to characterize the monotonicity of a function $f$ from the barcodes of $f$ and -$f$. Note that in Section 5.1, we characterize the monotonicity quantitatively using the barcode of the derivative.
@@
We infer the convexity of the function $m$ aforementioned in this subsection and compare the results with the respective SiZer maps. To infer the convexity of $m$ from the observed data, we need to consider the significant features in the observed barcode to apply step (i) of the proposed characterization. Note that, since we aim to infer the convexity of $m$ on the entire domain [-1, 1], the 0-dimensional feature with the largest persistence must be significant to proceed with step (ii), else we conclude that the observed structure is not convex (concave) significantly. The reason is that the features with shorter persistence represent structures in the subset of the domain [-1, 1] whereas here we need to infer the monotonicity of $m_{1}$ on [-1, 1]. For example, it can be seen in Table 3 that the feature with shorter persistence $[a_{2},b_{2}]$ is significant with the birth time $b_{2}<0$ (see, Figure 9). This implies that the observed structure is significantly non-increasing (see, Section 5.1) in the local regions determined by their respective critical points. This fact can also be validated by the SiZer maps in Figure 12 wherein the critical points $\boldsymbol{\hat{c}_{p}}$ corresponding to the observed features lie in red regions on the Sizer maps (see, Figure 12). This implies that the observed structure is significantly non-increasing in the local regions [-0.8, -0.6 ], [-0.6, -0.3], and [-0.5, -0.2], respectively, for $n=100,150,200$. Recall that green dots on the SiZer maps highlight the critical points $\boldsymbol{\hat{c}_{p}}$, and the observed features in the estimated barcodes =$\mathcal{B}=\{[a_{1},b_{1}],\ldots,[a_{p},b_{p}]\}$ are arranged in decreasing order of persistence.
We simulate noisy observations from $m$ using the noise distribution as the normal distribution with mean 0 and variance 0.1, truncated at [-1, 1]. We take the bandwidth $h=n^{-1/3}$ for $n\in\{100,150,200\}$. The true function (m), the estimated function ($\hat{m}$), the true derivative ($m_{1}$), and the estimated derivative ($\hat{m}_{1}$) are shown in Figure 9. It is evident from the visual inspection of the graph of $m_{1}$ in Figure 9 that $m_{1}$ is not a non-decreasing function implying that $m$ is not convex. We aim to infer this fact from the observed data using the proposed characterization.
Figure 9: An illustrative example for investigating the convexity of a regression function.
We estimate $m$ and $m_{1}$ using the normal and Cauchy kernel supported on [-1, 1]. We compute barcodes using the procedure given in Section 3 for $\hat{m}_{1}$ and -$\hat{m}_{1}$ shown in Figure 10 and 11 for the normal and Cauchy kernel, respectively. It can be seen in Figure 10 and 11 that there are two 0-dimensional features in the observed barcode of $m_{1}$ and -$m_{1}$ for $h=n^{-1/3},n\in\{100,150,200\}$. We need to assess the statistical significance of the observed features in the barcode to apply step (i) to infer the convexity of $m$.
Figure 10: True and estimated barcodes of $m_{1}$ and -$m_{1}$ for $n$ = 100, 150, 200.
@@
Now, we assess the statistical significance of these observed 0-dimensional features which is tabulated in Table 3, 4, 5, 6. It is evident from these tables that the observed 0-dimensional with the largest persistence is not significant for both $\hat{m}_{1}$ and -$\hat{m}_{1}$. Therefore, we infer that the observed structure is not convex. The SiZer maps in Figure 12 also suggest that the observed structure is not convex. Since there is a purple region between the red and blue region for the bandwidth $h=n^{-1/3}$ highlighted by the black dotted line. This suggests that the underlying function $m$ is not convex.
Table 3: Significance of convexity for Gaussian Kernel.
Barcode of $\boldsymbol{\hat{m}_{1}}$
$n=100$ $n=150$ $n=200$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$
$\mathbf{\hat{c}_{p}}$0.6 0.2 -0.6 -0.8 -0.6 0.8 -0.6 -0.3 -0.5 0.1 -0.5 -0.2
$\boldsymbol{\hat{\beta}_{1}}$3.5 0.8 -3.5 -1.3 -3.6 0.5 -3.6 -2.6 -3.6 0.3 -3.6 -2.2
$\mathbf{p}$-value $\boldsymbol{<10^{-14}}$ 0.1 $\boldsymbol{<10^{-14}}$ 0.007 $\boldsymbol{<10^{-16}}$ 0.2 $\boldsymbol{<10^{-16}}$ $\boldsymbol{<10^{-9}}$ $\boldsymbol{<10^{-16}}$ 0.4 $\boldsymbol{<10^{-16}}$ $\boldsymbol{<10^{-6}}$
@@
Table 4: Significance of convexity for -$m_{1}$.
Barcode of -$\boldsymbol{\hat{m}_{1}}$
$n=100$ $n=150$ $n=200$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$
$\mathbf{\hat{c}_{p}}$ 0.4 0.8 0.4 0.7 0.5 -0.8 0.5 0.7 0.7 -0.8 0.7 0.8
$\boldsymbol{\hat{\beta}_{1}}$1.4 0.6 -1.4 -0.6 -1.6 1.7 -1.6 -1.1 -1.2 1.9 -1.2 -1.1
$\mathbf{p}$-value 0.003 0.2 0.003 0.2 0.0005 0.0003 0.0005 0.01 0.003 $<10^{-5}$ 0.003 0.01
Table 5: Significance of convexity for Cauchy Kernel.
Barcode of $\boldsymbol{\hat{m}_{1}}$
$n=100$ $n=150$ $n=200$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$
$\mathbf{\hat{c}_{p}}$0.7 0.4 -0.5 -0.1 -0.7 0.4 -0.7 -0.07 -0.6 0.5 -0.5 -0.02
$\boldsymbol{\hat{\beta}_{1}}$2.2 1.3 -3.1 -0.9 -2.7 1.4 -3.02 -0.5 -3.5 1.7 -3.7 -0.68
$\mathbf{p}$-value $\boldsymbol{<10^{-6}}$ 0.02 $\boldsymbol{<10^{-9}}$ 0.1 $\boldsymbol{<10^{-8}}$ 0.01 $\boldsymbol{<10^{-9}}$ 0.3 $\boldsymbol{<10^{-13}}$ 0.003 $\boldsymbol{<10^{-13}}$ 0.2
Table 6: Significance of convexity for Cauchy Kernel for -$m_{1}$.
Barcode of -$\boldsymbol{\hat{m}_{1}}$
$n=100$ $n=150$ $n=200$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$
$\mathbf{\hat{c}_{p}}$ 0.4 -0.5 0.4 0.8 0.4 -0.5 0.7 0.9 0.7 -0.8 0.8 0.1
$\boldsymbol{\hat{\beta}_{1}}$1.3 3.3 -1.3 -0.4 -2.003 3.3 -1.1 -0.2 -1.3 0.9 -1.1 -0.2
$\mathbf{p}$-value 0.01 $<10^{-9}$ 0.01 0.4 0.0004 $<10^{-10}$ 0.02 0.7 0.005 0.05 0.02 0.7

### v4

@@ 5 Simulation studies
This section presents some applications of Theorem 3.1 to infer the monotonicity, convexity, and modality of smooth regression curves underlying the data. We simulate noisy observations from smooth regression curves and measure the statistical significance of the observed local structures using the significance measure proposed in Section 4.
In addition, we compare the results with a graphical device called SiZer map, an existing procedure for exploring structures such as peaks and valleys in smooth regression curves proposed by [6]. The SiZer map visualizes significant features such as bumps simultaneously over both location and scale (bandwidth) by using a color map. The idea is to highlight the regions in the location-scale space where the curve significantly increases and decreases. The idea is motivated by the fact that the significant bumps will be at zero crossings of the derivative between regions of significant increase and decrease. Hence, the name “SiZer” is given for “SIgnificant ZERo crossings of derivatives”. The color scheme on the SiZer map is as follows: Blue is used to indicate regions where the curve is significantly increasing; Red is used to indicate regions where the curve is significantly decreasing; Purple is used to indicate regions where the curve cannot be concluded to be either significantly increasing or decreasing; Gray is used to indicate regions where there are not enough data points to make statements about the significance of the features. The dotted curves in SiZer maps show effective window widths for each bandwidth as intervals representing $\pm 2h$.
In this section, we simulate noisy observations with the noise distribution as the truncated normal distribution supported on [-1, 1] with mean 0 and standard deviation 0.01. We estimate the regression function using the truncated standard normal and truncated standard Cauchy kernel supported on [-1, 1]. However, we compare the results with the SiZer maps only for the Gaussian kernel, since the conventional SiZer is applicable only for the Gaussian kernel.
@@ 5.1 Monotonicity
We characterize the monotonicity of a smooth and bounded regression function using the 0-dimensional features of the barcode of the first derivative. The main idea is that the birth and death times corresponding to the 0-dimensional feature with the largest persistence in a barcode provide the maximum and minimum value of the function, respectively (see, Figure 1). Therefore, we can extract the 0-dimensional feature with the largest persistence from the barcode of the first derivative to characterize the monotonicity of the function. In particular, if the death time of the feature is non-negative, that is, the minimum value of the first derivative is non-negative, then the function is non-decreasing. Similarly, if the birth time of the feature is non-positive, that is, the maximum value of the first derivative is non-positive, then the function is non-increasing.
Consider the regression function ${m(x)=e^{x},x\in}$ [-1, 1] to investigate its monotonicity from the kernel estimates of its derivative $m_{1}$. We obtain barcodes for the estimated derivative, $\hat{m}_{1}$, using the procedure mentioned in Section 3. The barcodes for the truncated standard Gaussian kernel are $\boldsymbol{\{[0.13,0.14]\}}$, $\boldsymbol{\{[0.15,0.17]\}}$ and $\boldsymbol{\{[0.17,0.20]\}}$ for the bandwidths $h=1.5(n=200),h=1.37(n=400)\text{ and }h=1.29(n=600)$, respectively. These barcodes are obtained for the filtration parameter $\epsilon=0.001$ and the bandwidths are chosen according to Equation (4.4) to assess the statistical significance of observed barcodes at $\alpha=0.05$ level of significance. Hence, by using Equation (4.3) for $z_{n,\alpha}=0.001$, we conclude that each of the observed bars are significant for the Gaussian kernel. Since the length of each bar (persistence) is greater than $5\sqrt{2}z_{n,\alpha}=\boldsymbol{0.007}$. This implies that the observed data suggest, with the 95% level of confidence, that the underlying function is strictly increasing, as the death times are positive for each of the bars.
The same conclusion also holds for the truncated standard Cauchy kernel. Since, the barcodes for the Cauchy kernel are $\boldsymbol{\{[0.07,0.13]\}}$, $\boldsymbol{\{[0.08,0.16]\}}$ and $\boldsymbol{\{[0.08,0.18]\}}$ for the bandwidths $h=2(n=200),h=1.8(n=400)\text{ and }h=1.7(n=600)$, respectively. Therefore, by Equation (4.3), all the observed bars are significant. Note that, here also, we have taken $\epsilon=0.001$ and the bandwidths are obtained using Equation (4.4) for $\alpha=0.05$. The estimated barcodes for both kernels are shown in Figure 2.
Figure 2: Estimated barcodes of $m_{1}$ for the Gaussian and Cauchy kernel truncated on [-1, 1].
Now, we compare the inferred structure from the barcode of $\hat{m}_{1}$ with the SiZer map of the data. It is evident from the SiZer map shown in Figure 3 that the observed structure is strictly increasing for each of the bandwidths used for the Gaussian kernel. Since the color of the scale-space region for each of the bandwidths (highlighted by the black dotted lines) is blue.
@@
Note that we do not use point-wise or uniform consistency of the estimators of $m$ or $m_{1}$ here. It is evident from Figure 4 and Figure 5 that neither $m$ nor $m_{1}$ is well estimated for each of the chosen bandwidths. Furthermore, the monotonicity (non-decreasingness) of the underlying regression function $m$ is not evident from the visual inspection of the estimates of $m$. However, we have shown that the local structures in $m$ can be inferred using the barcodes of $m_{1}$ without using pointwise or uniform consistency of $m$ or $m_{1}$. This is because Theorem 3.1 establishes the consistency of the estimated persistent homology of $m_{1}$ without stability results (see Theorem 2.1).
Figure 4: Kernel estimates of $m$ and $m_{1}$ for $h=1.5(n=200),1.37(n=400)\text{ and }1.29(n=600)$, using the truncated Gaussian kernel on [-1, 1].
Figure 5: Kernel estimates of $m$ and $m_{1}$ for $h=2(n=200),1.8(n=400)\text{ and }1.7(n=600)$, using the truncated Cauchy kernel on [-1, 1].
@@ 5.2 Convexity
Note that to characterize the monotone structure of $m_{1}$ quantitatively, we need to consider the barcodes of the second derivative of $m$. However, qualitatively, the monotonicity of $m_{1}$ can be captured by comparing the barcodes of $m_{1}$ and -$m_{1}$, as also suggested by one of the reviewers. The rationale behind the argument is that for a monotone function $f$, the super-level set $D_{t}=\{x:f(x)\geq t\}$ of $f$ has only a single connected component, for every $t\in\mathbb{R}$. However, this is not a necessary condition since if $f$ is a concave shape function, then also $D_{t}$ will have only a single connected component for every $t\in\mathbb{R}$. Therefore, to rule out such situations we have to make use the property of monotone functions that if $f$ is monotone increasing (decreasing) on an interval $[a,b]$ (here we take a = -1, b =1), then -$f$ will be monotone decreasing (increasing) on $[a,b]$. This property allows us to characterize the monotonicity of a function $f$ from the barcodes of $f$ and -$f$. Note that in Section 5.1, we characterize the monotonicity quantitatively using the barcode of the derivative.
@@
We infer the convexity of the function $m$ aforementioned in this subsection and compare the results with the respective SiZer maps. To infer the convexity of $m$ from the observed data, we need to consider the significant features in the observed barcode to apply step (i) of the proposed characterization. Note that, since we aim to infer the convexity of $m$ on the entire domain [-1, 1], the 0-dimensional feature with the largest persistence must be significant to proceed with step (ii), else we conclude that the observed structure is not convex (concave) significantly. The reason is that the features with shorter persistence represent structures in the subset of the domain [-1, 1], whereas here we need to infer the monotonicity of $m_{1}$ on [-1, 1].
We simulate noisy observations from $m$ and estimate $m_{1}$ for the bandwidths chosen according to Equation (4.4) for $\alpha=0.05$, and obtain barcodes for $\epsilon=0.001$. The estimated barcodes of $m_{1}$ using the truncated Gaussian kernel are $\boldsymbol{\{[-0.27,0.08]\}}$, $\boldsymbol{\{[-0.27,0.08]\}}$ and $\boldsymbol{\{[-0.28,0.09]\}}$ for $h=0.57(n=700)$, $h=0.56(n=800)$ and $h=0.55(n=900)$, respectively. Next, we estimate barcodes of -$m_{1}$ to investigate the convexity of $m$ using step (i) of the proposed characterization. The estimated barcodes of -$m_{1}$ for the same aforementioned parameters are $\boldsymbol{\{[-0.13,0.37]\}}$, $\boldsymbol{\{[-0.14,0.38],[-0.128,-0.122]\}}$ and $\boldsymbol{\{[-0.15,0.40],[-0.13,-0.12]\}}$ for $h=0.46(n=700)$, $h=0.45(n=800)$ and $h=0.44(n=900)$, respectively. The true and estimated barcodes for $m$ and -$m_{1}$ are shown in Figure 6.
Figure 6: True and estimated barcodes of $m_{1}$ and -$m_{1}$ for Gaussian kernel.
It is evident from Figure 6 that as $n$ increases, the true features of $m$ and -$m_{1}$ are recovered. The statistical significance of the observed bars is evident from Equation (4.3) for $z_{n,\alpha}=0.001$, as the chosen bandwidths are the solution of Equation (4.4) for $\alpha=0.05$ and $\epsilon=0.001$. Note that the second bar in the barcode of -$m_{1}$ is not significant, as its persistence is less than 0.007. However, as we increase $n$ from 800 to 900, the barcode of -$m_{1}$ consists of two significant bars. Hence, we conclude that the observed data suggest that the underlying function $m$ is not convex with the 95% level of confidence.
Alternatively, we obtain barcodes of $m_{1}$ using the truncated Cauchy kernel to infer the convexity of $m$ from the observed data. The estimated barcodes of $m_{1}$ are $\boldsymbol{\{[-0.18,0.04],[-0.12,-0.10]\}}$, $\boldsymbol{\{[-0.19,0.04],[-0.12,-0.10]\}}$ and $\boldsymbol{\{[-0.19,0.04],[-0.12,-0.10]\}}$ for $h=0.77(n=700)$, $h=0.75(n=800)$ and $h=0.73(n=900)$ , respectively. Subsequently, to proceed with step (i) of the proposed characterization, we estimate barcodes of -$m_{1}$. The barcodes of -$m_{1}$ are $\boldsymbol{\{[-0.07,0.24],[-0.046,-0.042]\}}$, $\boldsymbol{\{[-0.08,0.25],[-0.05,-0.04]\}}$ and $\boldsymbol{\{[-0.08,0.26],[-0.05,-0.04]\}}$ for $h=0.62(n=700)$, $h=0.61(n=800)$ and $h=0.6(n=900)$, respectively. These barcodes are shown in Figure 7.
@@
It is evident from the significance measure articulated in Equation (4.3), that all the observed bars are significant except for the second bar in the observed barcode of -$m_{1}$ for $n=700$. Note that here the significance measure is used for the bandwidths that solve Equation (4.4) for $\alpha=0.05$ and $\epsilon=0.001$. Therefore, a bar with a length less than 0.007 is considered to be insignificant. Hence, by step (i), the observed data suggests that $m$ is not convex with the 95% level of confidence, as there are two significant 0-dimensional features in the barcode of both $m_{1}$ and -$m_{1}$ as $n$ increases.
Thus, Theorem 3.1 provides a valid inference on the observed structure using estimates of $m$ and $m_{1}$ that are pointwise not well estimated, as can be seen in Figure 8. Moreover, the visual inspection of the estimated curves in Figure 8 does not reveal the true structure underlying the data.
Figure 8: Estimated curves using the Gaussian kernel for $h=0.57(n=700)$, $h=0.56(n=800)$ and $h=0.55(n=900)$.
Now, we compare the results with the SiZer map of the data shown in Figure 9. The SiZer maps in Figure 9 suggest that the observed structure is convex for the bandwidths chosen according to Equation (4.4) for the Gaussian kernel. This is because the SiZer maps detect a significantly decreasing slope followed by a significantly increasing slope of the estimated function in the regions highlighted by the black dotted lines on the SiZer maps. Note that this is in contrast with the true structure of the underlying function $m$, as $m$ is not convex. However, the proposed methodology correctly infers the true structure of the underlying regression function $m$.